# Robust Variational Physics-Informed Neural Networks

### 11 Dec 2023 v2

@@ Abstract
Keywords: Robustness, Variational Physics-Informed Neural Networks, Petrov-Galerkin formulation, Riesz representation, Minimum Residual principle, a posteriori error estimation.
@@ 1 Introduction
A natural continuation to PINNs addressing the aforementioned limitation is the so-called Variational PINNs (VPINNs) [29]. Here, the authors introduce a variational formulation of the underlying PDE within a Petrov-Galerkin framework. The solution is then approximated by a (deep) NN, whereas the test functions belong to linear vector spaces. Finally, the authors define a variational loss function to minimize during the training process. However, this methodology was not as popular as its predecessor. The main reason might be that the approach is sensible to the choice of the basis test functions. That is, given a discrete test space, the method’s loss, stability, and robustness heavily depends upon the choice of the basis functions of the given test space. In [4], authors present an a posteriori error analysis for discretizing elliptic boundary-value problems with VPINNs employing piecewise polynomials for the test space. As the loss function in VPINNs is, in general, not robust with respect to the true error (for example, the loss function can tend to zero even if the true error does not), they provide an error estimator employing classical techniques in finite element analysis to obtain practical information on the quality of the approximation. They also study in [5] the importance of selecting appropriate quadrature rules and the degree of piecewise polynomial test functions to guarantee optimal convergence rates for VPINNs when performing mesh refinements. Finally, in [30], the same authors from VPINNs introduced hp-VPINNs. In the latter, they employ piecewise polynomials for testing and allow for $hp$-refinements via domain decomposition by selecting non-overlapping test functions. However, the aforementioned issue of selecting an appropriate loss function for each problem remains also in this approach.
@@
In this work, we revisit the initial work on VPINNs from Kharazmi et al. in [29], and we introduce a Robust version of Variational Physics-Informed Neural Networks (RVPINNs). As in VPINNs, we consider a Petrov-Galerkin formulation of the PDE, approximate the solution employing NNs, and select a finite-dimensional test space. Our goal is to minimize the residual in the discrete dual norm. Therefore, we select a single test function that is the Riesz representation of the weak residual over the selected discrete test space. We select an inner product and span the Riesz representative on a discrete basis so the resulting loss function will include the inverse of the Gram matrix corresponding to the selected inner product. We prove that the norm of such test function is efficient and reliable, i.e., the norm of the true error is bounded from below and above (up to an oscillation term) by the norm of the residual representative. Moreover, as we show in the numerical results, our strategy is insensitive (unlike VPINNs) to the choice of the discrete basis spanning the test space (assuming that the numerical integration and the inversion of the Gram matrix are sufficiently accurate). It is easy to see that minimizing the proposed loss functional is equivalent to minimizing the test norm of the Riesz representation of the residual, which is the idea behind classical MinRes methods. Summarizing, we provide a general mathematical framework to define robust loss functionals in VPINNs. Our strategy relies on two ideas: (a) the appropriate selection of the inner product in the test space, ensuring the stability of the variational formulation, and (b) the selection of a single test function, that is the Riesz representation of the weak residual.
@@
In particular, if we select an orthonormal discrete basis with respect to the inner product in the test space, the Gram matrix becomes the identity, and we recover the original definition of the loss functional in VPINNs. The difference with classical VPINNs is that our strategy is robust and independent of the choice of the basis functions. Other works employ similar ideas closely related to minimum residual methods [9, 7]. In [49], the authors minimize the dual norm of the weak residual for symmetric variational formulations in $H^{1}$ and rectangular domains via a spectral decomposition. In [48], the same authors extend this method to time-harmonic Maxwell equations in the context of symmetric $H(curl)$-formulations. Also, in [50], in the context of Petrov-Galerkin formulations, the authors minimize the dual norm of the residual based on the concept of optimal testing form [20]. They approximate both the solution of the variational problem and the corresponding optimal test functions employing NNs by solving two nested deep Ritz problems. More recently, in [3], the authors interpolate the neural networks onto finite element spaces to represent the (partial) unknowns. This strategy overcomes, in particular, the challenges related to the imposition of boundary conditions and approximations of the NN derivatives during training.
@@
Finally, apart from the usual limitations of DL-based technologies (optimizer, integration, etc.), the largest bottleneck of RVPINNs at the moment is that in certain configurations, we need to invert the Gram matrix corresponding to the selected basis and inner product on the test space. We will explore how to optimize this step in the future. However, there are particular cases where the inversion of the Gram matrix is trivial (as in the spectral case) or easy to compute. For example, if we consider a strong variational formulation, the test inner product is $L^{2}$, and selecting piecewise discontinuous polynomials for testing the Gram matrix becomes block diagonal. On the other hand, for parametric problems, the Gram matrix may remain the same for different values of the parameters. Therefore, the inversion can be done offline, being valid for a large class of problems. Parametric problems are essential to solve inverse problems [24], so RVPINNs could be of great interest in this area.
@@
The article is organized as follows: In Section 2, we introduce the variational formulation of the model problem we consider in this article, the NN framework for approximation of the solution, and a brief overview of VPINNs. We also provide an alternative definition of VPINNs employing a single test function in the loss. Section 3 presents the methodology of RVPINNs and the connection with other methods based on VPINNs. Section 4 is devoted to the derivation of robust error estimates. We first prove that the residual representative is a reliable and efficient a posteriori error estimator of the true error in the sense of equivalent classes. Then, we demonstrate under the assumption of a local Fortin operator’s existence that the true error is equivalent to the residual error estimator up to an oscillation term. In Section 5 we test our method in several 1D and 2D advection-diffusion problems, showing the robustness of the approach. Finally, Section 6 summarizes the conclusions and future research lines.
@@ 2.1 Abstract framework
Problem (1) admits a unique solution by means of the Banach-Nečas-Babǔska Theorem, and the following a priori estimate holds (see [21, Theorem 1.1]):
@@
Additionally, it is well-known that (3) and (4) imply that the following adjoint inf-sup condition holds (see, e.g., [17, Theorem 1]:
@@ 2.2 Neural Network framework
Finally, for a given discrete space $V_{M}\subseteq V$, we consider the following Petrov-Galerkin discretization of (1):
@@ 2.3 Variational Physics-Informed Neural Networks
Variational Physics Informed Neural Networks (VPINNs) approximate the solution $u$ of problem (1) by minimizing a loss functional defined in terms of the weak residual (see Remark 1), and a discrete space $V_{M}\subseteq V$ with basis $\{\varphi_{m}\}_{m=1}^{M}$. That is, it approximates $u$ by solving the following minimization problem:
@@
with $\mathcal{L}_{r}\left(u_{\theta}\right)$ being a loss functional defined in terms of the weak residual and the discrete basis. This approach allows to consider scenarios where the solution belongs to less regular spaces than in collocation PINNs. For example, elliptic problems with solutions in $H^{2-\epsilon}(\Omega)$, with $0<\epsilon\leq 1$.
@@
where $C(\cdot)$ is a $V_{M}$-independent quadratic functional employed to impose boundary conditions for $u_{NN}$, that can be taken as $C(u_{\theta})=0$ if boundary conditions are strongly imposed in the DNN structure.
@@
If we assume sufficient approximability of the DNN structure and an ideal optimizer (solver), a minimizer for (13) for which the loss functional vanishes is also a solution to the Petrov-Galerkin problem (11). However, adopting such a loss functional may be computationally inefficient. Indeed, given the test space $V_{M}$, Eq. (13) depends heavily upon the choice of its basis. In particular, a simple re-scaling of one basis function (e.g., by making it sufficiently large) may easily lead to catastrophic results (cf. [49]). Consequently, there is a natural demand for robust loss definitions in practical VPINNs.
@@ 2.4 An alternative definition of VPINNs
The goal now is to define the loss functional as  (16) in terms of a particular discrete test function $\widetilde{\varphi}\in V_{M}$ such that VPINNs becomes robust. In the following Section, we present a strategy based on defining such a discrete function as the Riesz residual representative in $V_{M}$ with respect to the norm inducing the inf-sup stability (3).
@@ 3 Robust Variational Physics-Informed Neural Networks
In Robust Variational Physics-Informed Neural Networks (RVPINNs), we first introduce an inner product $(\cdot,\cdot)_{V}$111i.e., $||v||^{2}_{V}:=(v,v)_{V}$, for all $v\in V$., whose associated norm satisfies the inf-sup stability condition (3). Next, for a given trainable parameter $\theta$, we compute $\phi:=\phi(\theta)\in V_{M}$ being the solution of the following Galerkin problem:
@@
where $C(\cdot)$ is a $V_{M}$-independent quadratic functional to impose boundary conditions as in (13). Finally, we obtain the approximation of the solution of problem (1) by solving the following minimization problem (cf.,  [29]):
@@
In Section 4.2, we prove that $\|\phi\|_{V}$ is, up to an oscillation term, a local robust error estimation for the error $\|u-u_{\theta}\|_{U}$ under the assumption of the existence of a local Fortin’s operator.
@@ 4 Error estimates for RVPINNs
One of the main complexities for proving the robustness of the residual estimator $\|\phi\|_{V}$ is that the solution of the Petrov-Galerkin problem (11) may not have a solution or, if there exists, it may be non-unique since the space of all possible realizations for the NN structure defines a manifold instead of a finite-dimensional space (see, e.g., [5, Section 6.3]). Thus, standard FEM arguments based on a discrete inf-sup condition cannot be applied in this context.
@@ 4.1 A posteriori error estimates for RVPINNs in the sense of equivalence classes
Let $u\in U$ be the solution of the continuous problem (1); $u_{\theta}\in U_{NN}$ denote a DNN structure with trainable parameters $\theta\in\mathbb{R}^{S}$; $V_{M}\subseteq V$ denote a finite dimensional space, dotted with norm $\|\cdot\|_{V}$; and $\phi\in V_{M}$ be the solution of problem (17). It holds:
@@
Inequalities (33) show that, for any $\theta$, $\|\phi\|_{V}$ always, up to some multiplicative constants, defines an efficient (lower) and a reliable (upper) bound (thus, robust) estimation for the error $\|[u-u_{\theta}]\|_{U/U_{M}^{0}}$. This, in practice, implies that to minimize $\|\phi\|_{V}$ is equivalent to minimize $\|[u-u_{\theta}]\|_{U/U_{M}^{0}}$. Moreover, if we assume that problem (11) admits a solution, when $\|\phi\|_{V}\rightarrow 0^{+}$, $u-u_{\theta}$ converges to a function that belongs to $U_{M}^{0}$ as a consequence of the robustness. Thus, since $u\not\in U_{M}^{0}$, we conclude that $u_{\theta}$ converges to a function $u_{\theta^{\ast}}$ satisfying:
@@
Therefore, $u_{\theta}$ converges to a solution of the Petrov-Galerkin problem (11), as expected.
@@ 4.2 Energy norm error estimates based on local semi-discrete inf-sup condition
If we assume that the NN structure allows for solutions of the Petrov-Galerkin problem (11), we can also obtain a local reliable bound for $\|u-u_{\theta}\|_{U}$ through the following Assumption:
Assumption 1 (Local Fortin’s condition).
@@
since $\Pi_{\theta}v\in V_{M}$. Thus, in such a case, the residual representative $\|\phi\|_{V}$ is only a reliable estimate in the spirit of Theorem 2. Nevertheless, it is expected that, in the minimization procedure, the NN solution approaches a solution of the Petrov-Galerkin problem (11) (see Remark 4), implying that the NN solution will not belong to the space $U_{M}^{0}$ in a further iteration. Therefore, it is meaningful to assume that, after a sufficient number of iterations, the nonlinear solver will reach a parameter $\theta$ belonging to a neighborhood of a parameter $\theta^{\ast}$, in which $\theta^{\ast}$ is the local minimum, implying that a local semi-discrete inf-sup condition is satisfied, as stated in the following Lemma (cf. [23])
@@
$\dfrac{\alpha}{C_{\Pi}}\|w_{NN}\|_{U}\leq\sup_{0\neq v_{M}\in V_{M}}\dfrac{a(w_{NN},v_{M})}{\|v_{M}\|_{V}},\quad\forall\,w_{NN}\in U_{NN}^{\theta^{\ast}}:=\left\{u_{\theta}:\theta\in B(\theta^{\ast},R)\right\}.$ (36)
@@
$\sup_{0\neq v_{M}\in V_{M}}\dfrac{a(w_{NN},v_{M})}{\|v_{M}\|_{V}}\geq\sup_{0\neq v\in V}\dfrac{a(w_{NN},\Pi_{\theta}v)}{\|\Pi_{\theta}v\|_{V}}=\sup_{0\neq v\in V}\dfrac{a(w_{NN},v)}{\|\Pi_{\theta}v\|_{V}}\geq\dfrac{1}{C_{\Pi}}\sup_{0\neq v\in V}\dfrac{a(w_{NN},v)}{\|v\|_{V}}\geq\dfrac{\alpha}{C_{\Pi}}\|w_{NN}\|_{U}.$
@@
Proposition 6 (Local a priori error estimate).
@@
$\|u-u_{\theta^{\ast}}\|_{U}\leq\left(1+\dfrac{\mu C_{\Pi}}{\alpha}\right)\inf_{w_{NN}\in U_{NN}^{\theta^{\ast}}}\|u-w_{NN}\|_{U}.$ (37)
@@
First notice that, for all $w_{NN}\in U_{NN}^{\theta^{\ast}}$, it holds
$\dfrac{\alpha}{C_{\Pi}}\|u_{\theta^{\ast}}-w_{NN}\|_{U}\leq\sup_{0\neq v_{M}\in V_{M}}\dfrac{a(u_{\theta^{\ast}}-w_{NN},v_{M})}{\|v_{M}\|_{V}}=\sup_{0\neq v_{M}\in V_{M}}\dfrac{a(u-w_{NN},v_{M})}{\|v_{M}\|_{V}}\leq\mu\|u-w_{NN}\|_{U}.$
@@
$\|u-u_{\theta^{\ast}}\|_{U}\leq\|u-w_{NN}\|_{U}+\|u_{\theta^{\ast}}-w_{NN}\|_{U}.$
@@
Following [22], it can be proved that the converse of Lemma 5 is also valid. This could be useful in scenarios where proving the local semi-discrete inf-sup condition is simpler than proving the existence of a Fortin operator.

### 05 Mar 2024 v3

@@ Abstract
Keywords: Robustness, Variational Physics-Informed Neural Networks, Petrov-Galerkin formulation, Riesz representation, Minimum Residual principle, a posteriori error estimation, quasi-minimizers.
@@ 1 Introduction
A natural continuation to PINNs addressing the aforementioned limitation is the so-called Variational PINNs (VPINNs) [32]. Here, the authors introduce a variational formulation of the underlying PDE within a Petrov-Galerkin framework. The solution is then approximated by a (deep) NN, whereas the test functions belong to linear vector spaces. Finally, the authors define a variational loss function to minimize during the training process. However, this methodology was not as popular as its predecessor. The main reason might be that the approach is sensible to the choice of the basis test functions. That is, given a discrete test space, the method’s loss, stability, and robustness heavily depends upon the choice of the basis functions of the given test space. In [5], authors present an a posteriori error analysis for discretizing elliptic boundary-value problems with VPINNs employing piecewise polynomials for the test space. As the loss function in VPINNs is, in general, not robust with respect to the true error (for example, the loss function can tend to zero even if the true error does not), they provide an error estimator employing classical techniques in finite element analysis to obtain practical information on the quality of the approximation. They also study in [6] the importance of employing appropriate quadrature rules and the degree of piecewise polynomial test functions to guarantee optimal convergence rates for VPINNs when performing mesh refinements. Finally, in [33], the same authors from VPINNs introduced hp-VPINNs. In the latter, they employ piecewise polynomials for testing and allow for $hp$-refinements via domain decomposition by selecting non-overlapping test functions. However, the aforementioned issue of adopting an appropriate loss function for each problem remains also in this approach.
@@
In this work, we revisit the initial work on VPINNs from Kharazmi et al. in [32], and we introduce a Robust version of Variational Physics-Informed Neural Networks (RVPINNs). As in VPINNs, we consider a Petrov-Galerkin formulation of the PDE, approximate the solution employing NNs, and introduce a finite-dimensional test space. Our goal is to minimize the residual in the discrete dual norm. Therefore, we introduce a single test function that is the Riesz representation of the weak residual over the discrete test space. We define an inner product and span the Riesz representative on a discrete basis so the resulting loss function will include the inverse of the Gram matrix corresponding to the selected inner product. We prove that the norm of such test function is efficient and reliable, i.e., the norm of the true error is bounded from below and above (up to an oscillation term) by the norm of the residual representative. Moreover, as we show in the numerical results, our strategy is insensitive (unlike VPINNs) to the choice of the discrete basis spanning the test space (assuming that the numerical integration and the inversion of the Gram matrix are sufficiently accurate). It is easy to see that minimizing the proposed loss functional is equivalent to minimizing the test norm of the Riesz representation of the residual, which is the idea behind classical MinRes methods. Summarizing, we provide a general mathematical framework to define robust loss functionals in VPINNs. Our strategy relies on two ideas: (a) the appropriate selection of the inner product in the test space, ensuring the stability of the variational formulation, and (b) the adoption of a single test function, that is the Riesz representation of the weak residual.
@@
In particular, if we consider an orthonormal discrete basis with respect to the inner product in the test space, the Gram matrix becomes the identity, and we recover the original definition of the loss functional in VPINNs. The difference with classical VPINNs is that our strategy is robust and independent of the choice of the basis functions. Other works employ similar ideas closely related to minimum residual methods [12, 9]. In [2], the authors approximate the solution of symmetric and positive definite problems employing the discrete weak residual of the variational problem. For that, they define an adaptive construction of a sequence of finite-dimensional subspaces whose basis functions are realizations of a sequence of neural networks. In [54], the authors minimize the dual norm of the weak residual for symmetric variational formulations in $H^{1}$ and rectangular domains via a spectral decomposition. In [53], the same authors extend this method to time-harmonic Maxwell equations in the context of symmetric $H(curl)$-formulations. Also, in [55], in the context of Petrov-Galerkin formulations, the authors minimize the dual norm of the residual based on the concept of optimal testing form [23]. They approximate both the solution of the variational problem and the corresponding optimal test functions employing NNs by solving two nested deep Ritz problems. More recently, in [4], the authors interpolate the neural networks onto finite element spaces to represent the (partial) unknowns. This strategy overcomes, in particular, the challenges related to the imposition of boundary conditions and approximations of the NN derivatives during training.
@@
Finally, apart from the usual limitations of DL-based technologies (optimizer, integration, etc.), the largest bottleneck of RVPINNs at the moment is that in certain configurations, we need to invert the Gram matrix corresponding to the basis and inner product of the test space. We will explore how to optimize this step in the future. However, there are particular cases where the inversion of the Gram matrix is trivial (as in the spectral case) or easy to compute. For example, if we consider a strong variational formulation, the test inner product is $L^{2}$, and employing piecewise discontinuous polynomials for testing the Gram matrix becomes block diagonal. On the other hand, for parametric problems, the Gram matrix may remain the same for different values of the parameters. Therefore, the inversion can be done offline, being valid for a large class of problems. Parametric problems are essential to solve inverse problems [27], so RVPINNs could be of great interest in this area.
@@
The article is organized as follows: In Section 2, we introduce the variational formulation of the model problem we consider in this article, the NN framework for approximation of the solution, and a brief overview of VPINNs. We also provide an alternative definition of VPINNs employing a single test function in the loss. Section 3 presents the methodology of RVPINNs and the connection with other methods based on VPINNs. Section 4 is devoted to the derivation of robust error estimates. We first prove that the residual representative is a reliable and efficient a posteriori error estimator of the true error in the sense of equivalent classes. Then, we demonstrate under the assumption of a local Fortin operator’s existence that the true error is equivalent to the residual error estimator up to an oscillation term. Finally, we derive an a priori error estimate in the sense of quasi-optimizers. In Section 5 we test our method in several 1D and 2D advection-diffusion problems, showing the robustness of the approach. Finally, Section 6 summarizes the conclusions and future research lines.
@@ 2.1 Abstract framework
Problem (1) admits a unique solution by means of the Banach-Nečas-Babǔska Theorem, and the following a priori estimate holds (see [24, Theorem 1.1]):
@@
Additionally, it is well-known that (3) and (4) imply that the following adjoint inf-sup condition holds (see, e.g., [20, Theorem 1]:
@@ 2.2 Neural Network framework
In the following, we assume $U_{NN}\neq\emptyset$. Finally, for a given discrete space $V_{M}\subseteq V$, we consider the following Petrov-Galerkin-type discretization of (1):
@@
Remark 2 (Existence of solutions for the Petrov-Galerkin-type discretization).
The non-emptiness for $U_{NN}$ is a necessary assumption to guarantee the existence of solutions to problem (11), as an inadequate DNN structure setup could generate a manifold $U_{NN}$ that is not contained in $U$. This is the case of a strong variational formulation set in $U\subset H^{2}(\Omega)$ (cf. Remark 11) with activation functions that are not sufficiently smooth (e.g., ReLU). However, we also notice that non-emptiness is not a sufficient condition to guarantee the existence of solutions, as they could be meaningful only in a dense sense (i.e., when $S\rightarrow+\infty$, with $S$ denoting the number of trainable parameters). For instance, when considering smooth activation functions, but solutions are not smooth (cf. Section 5.5). The last can be circumvented by introducing the concept of quasi-minimizers, as we will see in Section 4.3.
@@ 2.3 Variational Physics-Informed Neural Networks
Variational Physics Informed Neural Networks (VPINNs) approximate the solution $u$ of problem (1) by minimizing a loss functional defined in terms of the weak residual (see Remark 1), and a discrete space $V_{M}\subseteq V$ with basis $\{\varphi_{m}\}_{m=1}^{M}$. That is, it approximates $u$ by solving the following minimization problem [^1] Note that this problem may not admit solutions as a consequence of Remark 2:
@@
with $\mathcal{L}_{r}\left(u_{\theta}\right)$ being a loss functional defined in terms of the weak residual and the discrete basis. This approach allows to consider scenarios where the solution belongs to less regular spaces than in collocation PINNs. For example, elliptic problems with solutions in $H^{2-\epsilon}(\Omega)$, with $\Omega\subset\mathbb{R}^{2}$ and $0<\epsilon\leq 1$.
@@
where $C(\cdot)$ is a $V_{M}$-independent non-negative-valued cost functional employed to impose data constraints, including boundary conditions, for $u_{NN}$ that can be taken as $C(u_{\theta})=0$ if boundary conditions are strongly imposed in the DNN architecture, and no data for the interpolation is available.
@@
If we assume sufficient approximability of the DNN architecture and an ideal optimizer (solver), a minimizer for (13) for which the loss functional vanishes is also a solution to the Petrov-Galerkin-type problem (11). However, adopting such a loss functional may be computationally inefficient. Indeed, given the test space $V_{M}$, Eq. (13) depends heavily upon the choice of its basis. In particular, a simple re-scaling of one basis function (e.g., by making it sufficiently large) may easily lead to catastrophic results (cf. [54]). Consequently, there is a natural demand for robust loss definitions in practical VPINNs.
Remark 4 (Constraint assumption).
We assume $C(\cdot)$ in (13) as a given cost functional since the present work focuses on the contribution in terms of the first term in (13). The efficient implementation considering both contributions is a topic that requires further exploration.
@@ 2.4 An alternative definition of VPINNs
The goal now is to define the loss functional as  (16) in terms of a particular discrete test function $\widetilde{\varphi}\in V_{M}$, depending upon $u_{\theta}$, such that VPINNs becomes robust. In the following Section, we present a strategy based on defining such a discrete function as the Riesz residual representative in $V_{M}$ with respect to the norm inducing the inf-sup stability (3).
@@ 3 Robust Variational Physics-Informed Neural Networks
In Robust Variational Physics-Informed Neural Networks (RVPINNs), we first introduce an inner product $(\cdot,\cdot)_{V}$222i.e., $||v||^{2}_{V}:=(v,v)_{V}$, for all $v\in V$., whose associated norm satisfies the inf-sup stability condition (3). Next, for a given trainable parameter $\theta$, we compute $\phi:=\phi(\theta)\in V_{M}$ being the solution of the following Galerkin problem:
@@
where $C(\cdot)$ is a $V_{M}$-independent functional to impose constraints as in (13). Finally, we obtain the approximation of the solution of problem (1) by solving the following minimization problem (cf.,  [32]):
@@
when the DNN architecture allows for solutions for the associated Petrov-Galerkin-type problem (see Remark 2).
@@
Remark 6 (Gram matrix inversion).
A practical implementation of RVPINNs requires the computation of $r(u_{\theta},\phi)$ following (22). Computing explicitly $G^{-1}$ could be computationally expensive in some scenarios, for instance, when considering standard FEM discrete test spaces defined in terms of piece-wise polynomials. Efficient inversions of the system (21) may require iterative solvers and preconditioners (see, e.g., [52]).
In Section 4.2, we prove that $\|\phi\|_{V}$ is, up to the constraint term $C(u_{\theta})$ and an oscillation term, a local robust error estimation for the error $\|u-u_{\theta}\|_{U}$ under the assumption of the existence of a local Fortin’s operator.
@@ 4 Error estimates for RVPINNs
One of the main complexities for proving the robustness of the residual estimator $\|\phi\|_{V}$ is that the solution of the Petrov-Galerkin problem (11) may not have a solution or, if there exists, it may be non-unique since the space of all possible realizations of the NN architecture defines a manifold instead of a finite-dimensional space (see, e.g., [6, Section 6.3]). Thus, standard FEM arguments based on a discrete inf-sup condition cannot be applied in this context.
@@ 4.1 A posteriori error estimates for RVPINNs in the sense of equivalence classes
Let $u\in U$ be the solution of the continuous problem (1); $u_{\theta}\in U_{NN}$ denote a DNN architecture with trainable parameters $\theta\in\mathbb{R}^{S}$; $V_{M}\subseteq V$ denote a finite dimensional space with norm $\|\cdot\|_{V}$; and $\phi\in V_{M}$ be the solution of problem (17). It holds:
@@
Inequalities (33) show that, for any $\theta$, $\|\phi\|_{V}$ always, up to some multiplicative constants, defines an efficient (lower) and a reliable (upper) bound (thus, robust) estimation for the error $\|[u-u_{\theta}]\|_{U/U_{M}^{0}}$. This, in practice, implies that to minimize $\|\phi\|^{2}_{V}$ $+\,C(u_{\theta})$ is equivalent to minimize $\|[u-u_{\theta}]\|^{2}_{U/U_{M}^{0}}$ $+\,C(u_{\theta})$. Moreover, if we assume that problem (11) admits a solution and consider $C(u_{\theta})=0$, $\|\phi\|_{V}\rightarrow 0^{+}$ implies that $u-u_{\theta}$ converges to a function that belongs to $U_{M}^{0}$ as a consequence of the robustness. Thus, since $u\not\in U_{M}^{0}$, we conclude that $u_{\theta}$ converges to a function $u_{\theta^{\ast}}$ satisfying:
@@
Therefore, $u_{\theta}$ converges to a solution of the Petrov-Galerkin-type problem (11), as expected.
@@ 4.2 Energy norm error estimates based on local semi-discrete inf-sup condition
If we assume that the DNN architecture allows for solutions of the Petrov-Galerkin problem (11), we can also obtain a local reliable bound for $\|u-u_{\theta}\|_{U}$ through the following Assumption [^3] This Assumption is not standard; it is an adaptation of the classical Fortin’s condition, commonly used in the discrete inf-sup stability analysis of mixed FEM formulations (see [8]) and, more recently, in the analysis of residual minimization-based FEMs, such as the DPG method (see [45]).:
Assumption 1 (Local Fortin’s condition I).
@@
since $\Pi_{\theta}v\in V_{M}$. Thus, in such a case, the residual representative $\|\phi\|_{V}$ is only a reliable estimate in the spirit of Theorem 2. Nevertheless, it is expected that, in the minimization procedure, the NN solution approaches a solution of the Petrov-Galerkin problem (11) (see Remark 7), implying that the NN solution will not belong to the space $U_{M}^{0}$ in a further iteration. Therefore, it is meaningful to assume that, after a sufficient number of iterations, the nonlinear solver will reach a parameter $\theta$ belonging to a neighborhood of a parameter $\theta^{\ast}$, in which $\theta^{\ast}$ is the local minimum, implying that a local semi-discrete inf-sup condition is satisfied, as stated in the following Lemma (cf. [26]).
@@
$\dfrac{\alpha}{C_{\Pi}}\|{u_{\theta}}\|_{U}\leq\sup_{0\neq v_{M}\in V_{M}}\dfrac{a({u_{\theta}},v_{M})}{\|v_{M}\|_{V}},\quad\forall\,{u_{\theta}}\in U_{NN}^{\theta^{\ast}}:={\operatorname*{span}}\left\{u_{\theta}:\theta\in B(\theta^{\ast},R)\right\}.$ (36)
@@
$\sup_{0\neq v_{M}\in V_{M}}\dfrac{a({u_{\theta}},v_{M})}{\|v_{M}\|_{V}}\geq\sup_{0\neq v\in V}\dfrac{a({u_{\theta}},\Pi_{\theta}v)}{\|\Pi_{\theta}v\|_{V}}=\sup_{0\neq v\in V}\dfrac{a({u_{\theta}},v)}{\|\Pi_{\theta}v\|_{V}}\geq\dfrac{1}{C_{\Pi}}\sup_{0\neq v\in V}\dfrac{a({u_{\theta}},v)}{\|v\|_{V}}\geq\dfrac{\alpha}{C_{\Pi}}\|{u_{\theta}}\|_{U}.$
@@
Proposition 6 (Local a priori error estimate I).
@@
$\|u-u_{\theta^{\ast}}\|_{U}\leq\left(1+\dfrac{\mu C_{\Pi}}{\alpha}\right)\inf_{{u_{\theta}}\in U_{NN}^{\theta^{\ast}}}\|u-{u_{\theta}}\|_{U}.$ (37)
@@
First notice that, for all ${u_{\theta}}\in U_{NN}^{\theta^{\ast}}$, it holds
$\dfrac{\alpha}{C_{\Pi}}\|u_{\theta^{\ast}}-{u_{\theta}}\|_{U}\leq\sup_{0\neq v_{M}\in V_{M}}\dfrac{a(u_{\theta^{\ast}}-{u_{\theta}},v_{M})}{\|v_{M}\|_{V}}=\sup_{0\neq v_{M}\in V_{M}}\dfrac{a(u-{u_{\theta}},v_{M})}{\|v_{M}\|_{V}}\leq\mu\|u-{u_{\theta}}\|_{U}.$
@@
$\|u-u_{\theta^{\ast}}\|_{U}\leq\|u-{u_{\theta}}\|_{U}+\|u_{\theta^{\ast}}-{u_{\theta}}\|_{U}.$
@@
Following [25], it can be proved that the converse of Lemma 5 is also valid. This could be useful in scenarios where proving the local semi-discrete inf-sup condition is simpler than proving the existence of a Fortin operator. Note that the continuity constant of the Fortin operator is a measure of the loss of stability. Ideally, we would like to find a Fortin operator which is indeed an orthogonal projection, i.e., $C_{\Pi}=1$. However, the construction of a Fortin operator [45] depends upon the formulation, the spaces, and the discretization. The construction of local Fortin operators for formulations involving NNs are out of the scope of this article.